Find the value of n for which the equation (n - 1)^2 uxx - y^2n uyy = ny^2n-1 uy , is
(1) parabolic
(2) hyperbolic
1
Expert's answer
2020-03-13T12:31:01-0400

Find the value of "n" for which the equation "(n-1)^2u_{xx}-y^{2n}u_{yy}=ny^{2n-1}u_y" is

(1) parabolic

(2) hyperbolic

Second-order PDE


"A(x,y)u_{xx}+B(x,y)u_{xy}+C(x,y)u_{yy}=F(x,y,u,u_x,u_y)"

The type of second-order PDE at a point "(x_0,y_0)" depends on the sign of the discriminant defined as


"\Delta(x_0,y_0)=\begin{vmatrix}
 B & 2A \\
 2C & B
\end{vmatrix}"

"\Delta(x_0,y_0)=\begin{vmatrix}
 0 & 2(n-1)^2 \\
 2(-y^{2n}) & 0
\end{vmatrix}=4(n-1)^2y^{2n}"

(1) parabolic: "\Delta(x_0,y_0)=0=>4(n-1)^2y^{2n}=0=>n=1 \ or\ y=0"

"n=1 \ or\ y=0"


(2) hyperbolic: "\Delta(x_0,y_0)>0=>4(n-1)^2y^{2n}>0=>n\not=1, y\not=0"

"n\not=1, y\not=0"